A triangle has an area of 72 square feet and a height of 9 feet? Find the length of its base.

Answer by jim_thompson5910(35256)

A = (bh)/2

72 = (b*9)/2

72*2 = 9b

144 = 9b

144/9 = b

16 = b

b = 16


So the base is 16 feet long